OpenStudy (anonymous):

Why is the range of \(y = \sin^{-1}\theta\) in Quadrants I and IV? I thought that the principal values were in Quadrants I and III?

OpenStudy (lgbasallote):

but the positive and negative reference angles are in Quadrant I and IV yes?

OpenStudy (anonymous):

Yes? Not entirely sure if I understand the concept of this.

OpenStudy (lgbasallote):

what don't you understand?

OpenStudy (anonymous):

Well, I don't quite understand what the range comes from? I understand that the principal values are the values that make \(y = \sin^{-1}\theta\) a function, but I don't quite understand what the change of domain and range is? Since I don't seem to understand the refernece angles you mentioned above for this, could you explain that?

OpenStudy (lgbasallote):

now....im not entirely sure what you're asking

OpenStudy (anonymous):

Well, going back to the original question, I was confused on why the range would be in Quadrants I and IV. Would I be getting the range of which quadrants it's in based on the principal values? For instance, the principal values are \(-\frac π2 ≤ \theta ≤ - \frac π2\) Would it mean that the range is in quadrants I and IV because if \(\theta\) is within that range, it would lie in Quadrants I and IV?

OpenStudy (anonymous):

I mean \(-\frac π2 ≤ \theta ≤ \frac π2\)

OpenStudy (anonymous):